At what height above the surface of the earth , does the weight of an object decrease to 0.8 of the original






1
Expert's answer
2022-03-08T19:36:04-0500
"g=\dfrac{GM}{R^2},\\\space\\
g_n=0.8g=\dfrac{0.8GM}{R^2}=\dfrac{GM}{R^2_n},\\\space\\
R_n=\dfrac{R}{\sqrt{0.8}}=7127\text{ km},"


which corresponds to a height above the Earth's surface of


"H=R_n-R=752\text{ km}."